OpenStudy (anonymous):

Which ordered pair is on the graph of the equation 2x + 5y = 4? A. (0, 1) B. (7, –2) C. (0, –1) D. (3, –2)

OpenStudy (loser66):

Any idea?

OpenStudy (loser66):

why?

OpenStudy (loser66):

If it is A) then x = 0 , y =1, hence 2x +5y =2*0 +5*1 = 5 \(\neq 4\), how A is the right answer?

OpenStudy (loser66):

hey, don't keep silent. Just test one by one from the options. If A is not correct one, then test B) , C) D) until you get the correct one. Don't give up

OpenStudy (loser66):

Ok, study or not... it's up to you. :)

OpenStudy (anonymous):

B?

OpenStudy (godlovesme):

I agree with @loser66 (x,y) try by plugging each value or graph and see if its right :)

OpenStudy (loser66):

show me why?

OpenStudy (loser66):

Hint: imitate exactly what I did above to show that B is the correct one.

OpenStudy (anonymous):

2 x 7 = 14 5 x -2 = -10 14 + -10 = 4

OpenStudy (loser66):

perfect!! you got my medal.

OpenStudy (anonymous):

Thank you!

OpenStudy (loser66):

yw